# Mamba-FCS: Joint Spatio-Frequency Feature Fusion, Change-Guided Attention, and SeK Loss for Enhanced Semantic Change Detection in Remote Sensing

### v1

@@ III-B VSSM Backbone Encoder- VMamba
VMamba has 3 variants named VMamba-Small, VMamba-Tiny and VMamba-Base. In our implementation, we utilize the VMamba-Base variant, which outperforms other variants with $C_{1}=128,C_{2}=256,C_{3}=512,C_{4}=1024,L_{1}=2,L_{2}=2,L_{3}=15$ and $L_{4}=2$.
@@ III-C Joint Spatio-Frequency Feature Fusion Mechanism
As illustrated in Figure 3, we propose a novel fusion mechanism $\mathcal{F}_{fusion}$ to enhance Change Detection in bi-temporal images that jointly exploits spatial cues and frequency ques to highlight subtle scene variations. At each stage $i$, the feature maps from pre-change ($X_{i}^{T_{1}}$) and post-change ($X_{i}^{T_{2}}$) images are processed as follows,
@@ III-C3 Attention-Based Spatial-Frequency Fusion
To enhance informative features and suppress noise, $X_{i}^{\mathrm{reduced}}$ is refined using the Convolutional Block Attention Module (CBAM) [51], as illustrated in Figure 3. First, channel attention generates a one-dimensional gate vector, which, after sigmoid [52] activation, is multiplied element-wise with $X_{i}^{\mathrm{reduced}}$ to produce the channel-refined feature $\hat{X}_{i}^{\mathrm{reduced}}$. Subsequently, spatial attention creates a two-dimensional saliency mask, which is sigmoid activated and multiplied point-wise with $\hat{X}_{i}^{\mathrm{reduced}}$, resulting in the final fused representation $X_{i}^{\mathrm{fused}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$.
Figure 4: The Architecture of Binary-Change Decoder, that refines change predictions through a top-down pathway from Stage IV to I (left panel). At each stage, fused features $X^{\text{fused}}_{i}$ are processed by a VSS block and a CBAM-based upsampling unit (right panel). Intermediate change masks $\{CM_{i}\}_{i=1}^{4}$ are extracted to be used in CGA.
@@ III-D Binary Change Decoder
As illustrated in Figure 4, the CBAM-based upsampling unit employs three parallel convolutions with kernels $1{\times}1$, $3{\times}3$, and $5{\times}5$ to gather context at complementary spatial scales, a strategy shown to enlarge the effective receptive field and sharpen object boundaries during decoder up-sampling[53, 54, 55, 56]. Each branch output is batch-normalized, concatenated, and re-weighted using CBAM’s channel and spatial attention mechanisms, which adaptively suppresses background noise. The result is combined with a convolutional branch of kernel size $1\times 1$ and interpolated to align with the H, W, and C of the next stage, before being passed to the subsequent, finer stage.
@@ III-E Semantic Map Decoders
The Semantic Map Decoders follow a similar top-down, multi-scale refinement strategy as $\mathcal{F}_{\text{BCD}}$, to produce timestamp-specific semantic maps that are conditioned on the change cues generated by $\mathcal{F}_{\text{BCD}}$ as seen in figure 5. Two weight-independent decoders, $\mathcal{F}_{\text{SCD}}^{T_{1}}$ and $\mathcal{F}_{\text{SCD}}^{T_{2}}$, share an identical architecture, allowing each to learn class priors specialised to the pre- and post-change scenes, respectively (see Eq.3).
@@ III-F Loss Function
Our network is trained using a composite loss that integrates widely adopted cross-entropy loss[57] and a mean Intersection over Union ($\mathrm{mIoU}$) regulariser[58] , alongside our novel Separated Kappa (SeK)[31] -based loss, which explicitly rewards semantic consistency within detected change regions.
@@
To address the inherent class imbalance in change detection tasks—where unchanged pixels typically dominate the dataset—and to improve boundary delineation accuracy, we incorporate an mIoU regularizer [58]. The mIoU metric provides balanced evaluation across all classes by computing the intersection over union for each class independently, thereby preventing the model from being biased toward the majority class. This regularization term has been shown to significantly improve segmentation performance[59].
Let $Q\in\mathbb{R}^{N\times N}$ represent the confusion matrix derived from predicted and Ground Truth (GT) labels, where $q_{i,j}$ denotes the count of pixels classified as class $i$ by the model but belonging to GT class $j$, for $i,j\in\{1,\cdots,N\}$,with class 1 signifying no-change. The mIoU loss can be defined as,
@@
$\mathrm{IoU}_{1}=\frac{q_{11}}{\sum_{k=1}^{2}q_{k1}+\sum_{k=1}^{2}q_{1k}-q_{11}},$ (11)
$\mathrm{IoU}_{2}=\frac{q_{22}}{\sum_{i=1}^{2}\sum_{j=1}^{2}q_{ij}-q_{11}},$ (12)
@@ III-F1 Seperated Kappa
$\hat{\eta}=\frac{\displaystyle\sum_{j=1}^{N}\bigl{(}\sum_{i}q_{ij}\bigr{)}\bigl{(}\sum_{i}q_{ji}\bigr{)}}{\bigl{(}\sum_{i,j}q_{ij}-q_{11}\bigr{)}^{2}}$ (16)
III-F2 SeK-based Loss
@@ III-F3 Overall Objective
Figure 5: Semantic Map Decoder Architecture. The decoder refines change predictions via a top-down pathway from Stage IV to Stage I. At each stage, encoder feature $X^{T_{j}}_{i}$ is modulated by the ChangeMap $CM_{i}$ using a Change-Guided Attention module, processed by a VSS block, and upsampled via a CBAM-based upsampling unit to produce the semantic map $Y^{T_{j}}$ for timestamp $T_{j}$.
@@ IV-A Datasets
SECOND Dataset [31]: The SECOND dataset is a semantic change detection dataset consisting of 4,662 pairs of aerial images captured at a spatial resolution of 0.53 meters/pixel. Each image is of size 512 $\times$ 512 pixels. The dataset covers urban regions including Hangzhou, Chengdu, and Shanghai, with annotations across six key land cover classes: non-vegetated surfaces, trees, low vegetation, water bodies, buildings, and playgrounds. This diversity allows for comprehensive evaluation of semantic change detection methods across various urban dynamics. We adopt the standard split of 2,968 training pairs and 1,694 test pairs for our experiments.
Landsat-SCD Dataset [60] : The Landsat-SCD dataset comprises 2,425 original image pairs with a spatial resolution of 30 meters/pixel, captured over Tumushuke, Xinjiang, China, from 1990 to 2020. These images, each sized 416 $\times$ 416 pixels, are annotated for changes across four land cover classes—farmland, desert, buildings, and water bodies—associated with 10 specific semantic change types. The dataset is divided into training, validation, and testing sets with a ratio of 3:1:1, corresponding to 1,455, 485, and 485 samples, respectively.
@@ IV-B Implementation Details
Our architecture is implemented in PyTorch and trained using the AdamW[61] optimizer with a learning rate of $1\times 10^{-4}$, weight decay of $5\times 10^{-3}$, and batch size of 4. Training is conducted for 30,000 iterations on the SECOND dataset and 50,000 iterations on the Landsat-SCD dataset. During training, we apply comprehensive data augmentations including geometric transformations such as random rotations, horizontal and vertical flips, as well as photometric augmentations involving random adjustments to saturation, contrast, and brightness to enhance robustness against illumination and color variations. The complete source code, configuration files and pre-trained models will be publicly available at https://github.com/Buddhi19/Mamba-FCS.git upon publication.
@@ V Results and Discussion
This section presents the experimental results on benchmark datasets. We first perform qualitative and quantitative ablation studies to evaluate the impact of individual components in the proposed architecture on SECOND dataset. Next, we compare our method’s performance, both qualitatively and quantitatively, against state-of-the-art (SOTA) methods in SCD for both SECOND dataset and Landsat-SCD dataset.
TABLE I: Quantitative results of the ablation study.
Method CGA SeK $D_{i}$ $F_{i}^{T_{1}},F_{i}^{T_{2}}$ OA (%) $F_{scd}$ (%) mIoU (%) SeK (%)
w/o CGA 88.08 63.61 73.93 24.07
w/o SeK 88.43 65.09 73.67 24.71
w/o Difference Branch 88.34 65.11 73.77 24.83
w/o FFT2 Branch 87.86 64.52 73.16 24.03
Mamba-FCS 88.62 65.78 74.07 25.50
V-A Ablation Studies
Our proposed architecture, Mamba-FCS, comprises three key components, (i) Joint Spatio-Frequency Fusion Mechanism, (ii) Change-Guided Attention (CGA) Module, and (iii) SeK Loss. To evaluate their individual contributions, we conducted ablation studies on the SECOND dataset by systematically removing each component. Specifically, we assessed the model’s performance with and without CGA and with and without SeK Loss, both quantitatively and qualitatively. Within the fusion mechanism, we evaluated the FFT2 branch both quantitatively and qualitatively, and additionally the Difference branch quantitatively.
Quantitative results, are reported in Table I. Qualitative analyses for the FFT2 branch, CGA, and SeK Loss are presented in Figures 6, 7, and 8, respectively. Each figure is structured with four columns. The first displays input bi-temporal images, the second shows the Ground Truth (GT), the third presents the model’s output without the ablated component, and the fourth depicts the full model’s output. Red boxes highlight regions where each component significantly enhances performance. The results in Table I and the qualitative visualizations collectively demonstrate that each component plays a critical role in improving the model’s performance for SCD.

### v2

@@ III-B Visual State Space Model Based Backbone Encoder- VMamba
VMamba has 3 variants named VMamba-Small, VMamba-Tiny and VMamba-Base. In our implementation, we utilize the VMamba-Base variant, which outperforms other variants with configuration given in table I.
TABLE I: Configuration details for stages $i\in\{1,2,3,4\}$. $C_{i}$ denotes the channel depth, $L_{i}$ represents the number of VSS blocks in each encoder stage, and $H_{i}$ and $W_{i}$ denote the height and width relative to the original image dimensions $H$ and $W$.
Stage - $i$ 1 2 3 4
$C_{i}$ 128 256 512 1024
$L_{i}$ 2 2 15 2
$H_{i}$ $H/4$ $H/8$ $H/16$ $H/32$
$W_{i}$ $W/4$ $W/8$ $W/16$ $W/32$
@@ III-C Joint Spatio-Frequency Feature Fusion Mechanism
As illustrated in Figure 3, we propose a novel fusion mechanism $\mathcal{F}_{fusion}$ to enhance CD in bi-temporal images that jointly exploits spatial cues and frequency cues to highlight subtle scene variations. At each stage $i$ in the Binary Change Decoder, the feature maps from pre-change ($X_{i}^{T_{1}}$) and post-change ($X_{i}^{T_{2}}$) images are fused using the Fusion Block as illustrated in figure 4 to form $X_{i}^{\mathrm{fused}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ as follows,
Figure 4: (a) Architecture of the Binary Change Decoder for generating the Binary Change Map $Y_{BCD}$. At each stage, encoder features from two time points ($X_{i}^{T_{1}}$ and $X_{i}^{T_{2}}$) are fused through a fusion block to obtain $X^{\text{fused}}_{i}$. The fused features are then passed through a VSS block followed by a CBAM-based upsampling unit. Point-wise addition progressively integrates multi-scale information, while intermediate change maps $\{CM_{i}\}_{i=1}^{4}$ are extracted to support the Change-Guided Attention (CGA) module. (b) The architecture of the CBAM-based Upsampling Block, which reduces $C_{i}$ and increases the $H_{i}$ and $W_{i}$ for the next stage.
@@ III-C3 Attention-Based Spatial-Frequency Fusion
To enhance informative features and suppress noise, we refine the reduced feature map $X_{i}^{\mathrm{reduced}}$ using the Convolutional Block Attention Module (CBAM) [47], as illustrated in Fig. 3(b). CBAM first applies channel attention to produce $\mathbf{g}^{C}_{i}$ as,
$\mathbf{g}^{C}_{i}=\sigma\left(\mathrm{MLP}\big(\mathrm{GAP}(X_{i}^{\mathrm{reduced}})\big)+\mathrm{MLP}\big(\mathrm{GMP}(X_{i}^{\mathrm{reduced}})\big)\right),$ (8)
where $\sigma$ is the sigmoid activation, $\mathrm{GAP}$ and $\mathrm{GMP}$ are global average and max pooling over spatial dimensions, and the two pooled vectors are passed through a shared bottleneck MLP [47]. The resulting channel gate is applied to $X_{i}^{\mathrm{reduced}}$ as
$\hat{X}_{i}^{\mathrm{reduced}}=\mathbf{g}^{C}_{i}\odot X_{i}^{\mathrm{reduced}},$ (9)
with $\odot$ denoting element-wise multiplication.
Next, spatial attention is computed on $\hat{X}_{i}^{\mathrm{reduced}}$:
[…]
where $\mathrm{Avg}_{c}$ and $\mathrm{Max}_{c}$ perform average and max pooling along the channel dimension, $f^{7\times 7}$ is a $7\times 7$ convolution, and denotes channel-wise concatenation. The final fused representation is
$X_{i}^{\mathrm{fused}}=\mathbf{g}^{S}_{i}\odot\hat{X}_{i}^{\mathrm{reduced}}.$ (11)
Intuitively, $\mathbf{g}^{C}_{i}$ reweights feature channels based on global context (blue bars in Fig. 3(b)), while $\mathbf{g}^{S}_{i}$ emphasizes salient spatial locations (blue squares), yielding the refined fused representation $X_{i}^{\mathrm{fused}}$.
Figure 5: (a) Architecture of the Semantic Map Decoder for the $j^{\text{th}}$ time stamp ($j\in\{1,2\}$). At each stage, the encoder feature $X_{i}^{T_{j}}$ is refined using the corresponding change map $CM_{i}$ through a Change-Guided Attention module, producing $\hat{X}_{i}^{T_{j}}$. Point-wise addition is employed at Stages I–III to progressively integrate multi-scale information. The refined features are then processed by a VSS block and upsampled via a CBAM-based upsampling unit. The final output is the semantic map $Y^{T_{j}}$ for timestamp $T_{j}$.
@@ III-D Binary Change Decoder
As illustrated in Figure 4(b), the CBAM-based upsampling unit employs three parallel convolutions with kernels $1{\times}1$, $3{\times}3$, and $5{\times}5$ to gather context at complementary spatial scales, a strategy shown to enlarge the effective receptive field and sharpen object boundaries during decoder up-sampling[8, 37, 46, 50]. Each branch output is batch-normalized, concatenated, and re-weighted using CBAM’s channel and spatial attention mechanisms as described in equations (8) – (11), which adaptively suppresses noise. The result is combined with a convolutional branch of kernel size $1\times 1$ and interpolated to align with the H, W, and C of the next stage, before being passed to the subsequent, finer stage.
@@ III-E Semantic Map Decoders
The Semantic Map Decoders follow a similar top-down, multi-scale refinement strategy as $\mathcal{F}_{\text{BCD}}$, to produce timestamp-specific semantic maps that are conditioned on the change cues generated by $\mathcal{F}_{\text{BCD}}$ as seen in figure 5. Two weight-independent decoders, $\mathcal{F}_{\text{SCD}}^{T_{1}}$ and $\mathcal{F}_{\text{SCD}}^{T_{2}}$, share an identical architecture, allowing each to learn class priors specialised to the pre and post-change scenes, respectively (see Eq.3).
@@ III-F Loss Function
Our network is trained using a composite loss that integrates widely adopted cross-entropy loss[33] and a mean Intersection over Union ($\mathrm{mIoU}$) regulariser[34] , alongside our novel Separated Kappa (SeK)[49] inspired loss, which explicitly rewards semantic consistency within detected change regions.
@@
To address the inherent class imbalance in change detection tasks, where unchanged pixels typically dominate the dataset, and to improve boundary delineation accuracy, we incorporate an mIoU regularizer [34]. The mIoU metric provides balanced evaluation across all classes by computing the intersection over union for each class independently, thereby preventing the model from being biased toward the majority class. This regularization term has been shown to significantly improve segmentation performance[59].
Figure 6: Illustration of the confusion matrix $Q$ and the regions used in computation of SeK. (a) Full confusion matrix $Q$ with entries $q_{i,j}$. (b) $\mathrm{IoU}_{2}$ – the sum of red cells represent the numerator (all predicted–change vs. true–change entries), and the sum of green cells represent the denominator (all entries excluding $q_{11}$). (c) $\hat{\rho}$ – the sum of red diagonal cells correspond to the numerator (correctly classified change classes), and the sum of green cells correspond to the denominator (all change-related entries, i.e. all except $q_{11}$). (d) $\hat{\eta}$ – the numerator is computed from sum of the products of row and column sums of the red region, while the square of the sum of green cells indicates the denominator.
As illustrated in figure 6(a), $Q\in\mathbb{R}^{N\times N}$ represent the confusion matrix derived from predicted and Ground Truth (GT) labels, where $q_{i,j}$ denotes the count of pixels classified as class $i$ by the model but belonging to GT class $j$, for $i,j\in\{1,\cdots,N\}$,with class 1 signifying no-change. We define mIoU loss as,
@@
$\mathrm{IoU}_{1}=\frac{q_{11}}{\sum_{k=1}^{N}q_{k1}+\sum_{k=1}^{N}q_{1k}-q_{11}},$ (15)
$\mathrm{IoU}_{2}=\frac{\sum_{i=2}^{N}\sum_{j=2}^{N}q_{ij}}{\sum_{i=1}^{N}\sum_{j=1}^{N}q_{ij}-q_{11}},$ (16)
@@ III-F1 Seperated Kappa
$\hat{\eta}\;=\;\frac{\displaystyle\sum_{j=2}^{N}\Big(\sum_{i=2}^{N}q_{ji}\Big)\Big(\sum_{i=2}^{N}q_{ij}\Big)}{\Big(\displaystyle\sum_{i=2}^{N}\sum_{j=2}^{N}q_{ij}\Big)^{2}}.$ (20)
As illustrated in figure  6(b), $\mathrm{IoU}_{2}$ quantifies the overlap between predicted and true change pixels. In figure  6(c), $\hat{\rho}$ measures semantic accuracy within change regions by focusing on the diagonal elements of the reduced confusion matrix. As shown in figure 6(d), $\hat{\eta}$ accounts for chance agreement within the same change-only matrix. Thus, SeK effectively emphasizes semantic consistency in change regions while mitigating the influence of the non-change class.
III-F2 SeK Inspired Loss
@@
Validity of SeK Inspired Loss
For training, we adopt a differentiable “soft confusion matrix” formulation of SeK. Let $\Omega$ denote the set of pixels and let $y_{x}\in\{1,\dots,N\}$ be the ground-truth change type (with $1$ denoting no-change) at pixel $x\in\Omega$. For each pixel, the network outputs logits $z(x)$ and softmax probabilities $p_{i}(x)=\mathrm{softmax}(z(x))_{i}$. The $(i,j)$-th entry of the soft confusion matrix $Q$ is then
$q_{ij}=\sum_{x\in\Omega}\mathbf{1}[y_{x}=j]\,p_{i}(x),$ (24)
where $\mathbf{1}[\cdot]$ is the indicator function. With this construction, all quantities entering the SeK expression (the $q_{ij}$, $\mathrm{IoU}_{2}$, $\hat{\rho}$, $\hat{\eta}$, $\mathrm{SeK}_{1}$, $\mathrm{SeK}_{2}$, and $\mathrm{SeK}_{\text{avg}}$) become compositions of sums, products, and ratios of softmax probabilities, and are therefore differentiable with respect to the network logits.
The final loss $\mathcal{L}_{\mathrm{SeK}}$ in (23) is obtained by applying a clipping operator and a negative logarithm to $\mathrm{SeK}_{\text{avg}}$. For $\mathrm{SeK}_{\text{avg}}>0$, we have $\overline{\mathrm{SeK}}=\mathrm{SeK}_{\text{avg}}$ and
$\frac{\partial\mathcal{L}_{\mathrm{SeK}}}{\partial\mathrm{SeK}_{\text{avg}}}=-\frac{1}{\mathrm{SeK}_{\text{avg}}+\varepsilon},$
which is finite for all $\mathrm{SeK}_{\text{avg}}>0$. The clipping at zero introduces a non-differentiable point at $\mathrm{SeK}_{\text{avg}}=0$, but this is a standard subgradient situation (analogous to ReLU), and in practice we observe $\mathrm{SeK}_{\text{avg}}>0$ during training. Since $\overline{\mathrm{SeK}}\in[0,1]$, the loss $\mathcal{L}_{\mathrm{SeK}}=-\log(\overline{\mathrm{SeK}}+\varepsilon)$ is bounded between two finite constants and decreases monotonically as semantic agreement in changed regions improves. Hence, the proposed SeK loss is well-defined, bounded from below, and (almost everywhere) differentiable with respect to the model parameters, and can be safely optimized via backpropagation.
@@ IV-A Datasets
SECOND Dataset [49]: The SECOND dataset is a semantic change detection dataset consisting of 4,662 pairs of aerial RGB images captured at a spatial resolution of 0.53 meters/pixel. Each image is of size 512 $\times$ 512 pixels. The dataset covers urban regions including Hangzhou, Chengdu, and Shanghai, with annotations across six key land cover classes: non-vegetated surfaces, trees, low vegetation, water bodies, buildings, and playgrounds. This diversity allows for comprehensive evaluation of semantic change detection methods across various urban dynamics. We adopt the standard split of 2,968 training pairs and 1,694 test pairs for our experiments.
Landsat-SCD Dataset [51] : The Landsat-SCD dataset comprises 2,425 original RGB image pairs with a spatial resolution of 30 meters/pixel, captured over Tumushuke, Xinjiang, China, from 1990 to 2020. These images, each sized 416 $\times$ 416 pixels, are annotated for changes across four land cover classes, namely farmland, desert, buildings, and water bodies, associated with 10 specific semantic change types. The dataset is divided into training, validation, and testing sets with a ratio of 3:1:1, corresponding to 1,455, 485, and 485 samples, respectively.
TABLE II: Quantitative results of the ablation study performed on the SECOND and Landsat-SCD datasets. Reported values are averaged over five independent runs.
SECOND Landsat-SCD
Method CGA SeK $F_{i}^{T_{1}},F_{i}^{T_{2}}$ OA (%) $F_{\text{scd}}$ (%) mIoU (%) SeK (%) OA (%) $F_{\text{scd}}$ (%) mIoU (%) SeK (%)
Base Model 86.08 62.49 72.65 23.70 95.17 85.54 86.48 52.50
w/o CGA 88.08 63.61 73.93 24.07 95.75 87.26 87.89 56.48
w/o SeK Loss 88.43 65.09 73.67 24.71 95.77 86.76 87.52 56.56
w/o FFT2 Branch 87.86 64.52 73.16 24.03 96.06 88.72 88.27 58.75
w/o SeK Loss, w/o CGA 87.26 63.30 72.89 23.42 95.68 87.58 87.30 55.96
Mamba-FCS 88.62 65.78 74.07 25.50 96.25 89.27 88.81 60.26
@@ IV-B Implementation Details
Our architecture is implemented in PyTorch and trained using the AdamW [30] optimizer with a learning rate of $1\times 10^{-4}$, weight decay of $5\times 10^{-3}$, and a batch size of 4. Training is conducted for 30,000 iterations on the SECOND dataset and 50,000 iterations on the Landsat-SCD dataset. To improve robustness against variations in illumination, viewpoint, and appearance, we apply extensive data augmentation, including random rotations, horizontal and vertical flips, and photometric adjustments to saturation, contrast, and brightness.
All experiments are performed on a workstation equipped with an NVIDIA RTX A6000 GPU (48 GB VRAM), a 32-core CPU, and 126 GB of system memory. Under these settings, training requires approximately 24 hours for SECOND and 36 hours for Landsat-SCD. On this hardware, Mamba-FCS records a peak GPU memory usage of 1152.65 MB and an average inference time of 284.69 ms per $512\times 512$ bi-temporal RGB pair. For reproducibility, we fix the random seed to 42 for Python and PyTorch (including the data loader) in all experiments. The complete source code, configuration files, and pre-trained models are publicly released at https://github.com/Buddhi19/Mamba-FCS.git upon publication.
@@ V Results and Discussion
This section presents the experimental results on benchmark datasets. We begin with quantitative ablation on the SECOND and Landsat-SCD datasets, and qualitative ablation on the SECOND dataset isolating the contributions of the Joint Spatio–Frequency Fusion mechanism, the Change-Guided Attention (CGA) module, and the SeK-inspired loss, as well as their combined effect. We then analyse the loss design itself by comparing the SeK-inspired term with alternative imbalance-aware losses and by studying the sensitivity of its weighting coefficients. Next, we benchmark Mamba-FCS against SOTA CNN, Transformer, and Mamba based SCD methods on both datasets, using both scalar metrics and qualitative visualisations, including dataset-level from–to confusion matrices and signed error maps to examine land-cover transition patterns. Finally, we report the computational cost and efficiency of Mamba-FCS relative to representative Transformer and Mamba baselines to assess its suitability for large-scale deployment.
V-A Ablation Studies on Key Contributions
Our proposed architecture, Mamba-FCS, comprises three key components: (i) a Joint Spatio–Frequency Fusion mechanism, (ii) a Change-Guided Attention (CGA) module, and (iii) a SeK-inspired loss. To isolate their individual contributions, we conduct ablation experiments on the SECOND and Landsat-SCD datasets by systematically removing each component. Specifically, we compare variants with and without CGA and with and without the SeK-inspired loss using both quantitative metrics and qualitative visualizations. Within the fusion mechanism, we further analyze the role of the FFT2 branch through combined quantitative and qualitative evaluation. Finally, we assess the joint impact of simultaneously disabling both the CGA module and the SeK-inspired loss, as well as removing all three components, to highlight their complementary benefits.